OpenStudy (anonymous):

medal anyone?? What are the real or imaginary solutions of the polynomial equation? 27x^3 + 125 = 0

OpenStudy (anonymous):

take the cubic root of both sides. 3x + 5 = 0 x = -5/3

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

oh snap didnt see you sorry lol and are you just trying to get x by itself?

OpenStudy (anonymous):

you can't do it like that... there may be (actually, there are) imaginary or complex solutions also.

OpenStudy (anonymous):

oh jeez.. sounds fun

OpenStudy (anonymous):

I'm terrible at factoring polynomials like this, but you can factor out (3x+5) from the polynomial, leaving: (3x+5)(9x^2 - 15x + 25)

OpenStudy (anonymous):

I'm not sure "fun" is the right word! The first part would lead to a real solution of x = -5/3 The second part has two complex solutions you can get from the quadratic formula.

OpenStudy (anonymous):

hahaha i got it now, i have another one too and if you can get it i'll follow and medal you for it

OpenStudy (anonymous):

Use the Remainder Theorem to find the remainder when P(x) = x^4 – 9x^3 – 5x^2 – 3x + 4 is divided by x + 3.

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

@cwrw238

OpenStudy (sirm3d):

The Remainder Theorem says that the remainder R is P(a). Write x+3 as x - (-3). So a = -3. Evaluate P(-3)